Suppose that a firm manufactures wallets. The firm's fixed cost is $102,972, and the variable cost is $3 per wallet. If each wallet sells for $71, how many wallets need to be sold in order for the firm's profit to be $54,104?
Answer #1

Suppose that a firm manufactures wallets. The firm's fixed cost is $102,972, and the variable cost is $3 per wallet. If each wallet sells for $71, how many wallets need to be sold in order for the firm's profit to be $54,104?

Revenue-Total fixed cost-Total variable cost=Profit
Now, suppose the total number of wallets sold be x.
Total variable cost=(Variable cost per wallet)*(Total number of wallets sold)=3x
Revenue=(Selling price per wallet)*(Total number of wallets sold)=71x

=>71x-102972-3x=54104
68x-102972=54104
68x=54104+102972
68x=157076
=>x=157076/68=2309.941176 or 2310 (Approximately)